OpenStudy (anonymous):

What is the simplified form of the expression the square root of the quantity 16 times c to the eighth power. - the square root of c squared. + 2the square root of c squared. + the square root of the quantity 4 times c squared.

OpenStudy (anonymous):

helppp please i have a test tomorrow and need help :(

OpenStudy (anonymous):

Welcome to OpenStudy! I love your display photo, okay let me look at the problem and try my best to help you!

terenzreignz (terenzreignz):

You're not trying hard enough :D LOL Just kidding @Kitt020912 :3

terenzreignz (terenzreignz):

The asker's gone? Hmmm... \[\Large \sqrt{16c^8}-\sqrt{c^2}+2\sqrt{c^2}+\sqrt{(4c)^2}\]

OpenStudy (anonymous):

Okay, that makes more sense now Cx

terenzreignz (terenzreignz):

I do love making sense :D